OpenStudy (anonymous):

3/3x + 1/x+4 = 10/7x solve for x

OpenStudy (anonymous):

-1/7

OpenStudy (anonymous):

\[\frac{3}{3x}+\frac{1}{x}+4=\frac{10}{7x}\] I'm assuming that's your expression. It's difficult to tell what you mean when you don't have parentheses. The first thing you need to do is to bring the variable up to the numerator. An easy way to do that is to multiply the equation through by x. \[x*\frac{3}{3x}+x*\frac{1}{x}+x*4=x*\frac{10}{7x}\] Since you multiplied everything by x, you are only scaling the expression by the variable x, and not changing it. The new expression should be: \[\frac{3}{3}+\frac{1}{}+4x=\frac{10}{7}\] Now, there is only 1 x value and you can simply add the fractions then isolate x to find its value.